OpenStudy (anonymous):

A rectangular equation that is not a function may be a function in the polar coordinate system. true or false? And please explain why.

OpenStudy (anonymous):

whats ur definition to a function :O and yes its true

OpenStudy (anonymous):

Consider the polar function \(r=1\), the unit circle centered at the origin. In Cartesian (rectangular) coordinates, the equation becomes \(x^2+y^2=1\), which is not a function because the equation is true for \((x,y)=\left(\dfrac{1}{\sqrt2},-\dfrac{1}{\sqrt2}\right)\) and \((x,y)=\left(\dfrac{1}{\sqrt2},\dfrac{1}{\sqrt2}\right)\). Graphically, you can think of this in terms of the vertical line test: